I need to select at least two odd numbers, two even numbers and zero for a formula that yields prime numbers...one such formula is x -x + 41. Select some numbers for x, substitute them in a formula and see if prime numbers occur. Try to find a number for x that when substituted in the formula yields a composite number
Answer by stanbon(75887)

I need to select at least two odd numbers, two even numbers and zero for a formula that yields prime numbers...one such formula is x -x + 41.
----------
Select some numbers for x, substitute them in a formula and see if prime numbers occur.
f(x) = x^2-x+41
odd: f(3) = 9-3+41 = 47 is prime
odd: f(7) = 49-7+41 = 83 is prime
----
even:f(4) = 16-4+41 = 53 is prime
even:f(6) = 36-6+41 = 71 is prime
-----
zero: f(0) = 41 is prime
----
Try to find a number for x that when substituted in the formula yields a composite number
f(41) = 1681 which is 41^2
-------------------
Cheers,
Stan H.